OpenStudy (anonymous):

Meredith works at a horse stable. Her job is to exercise the horses. The stable has a circular area that Meredith runs the horses within. The diameter of the fencing is 16 yards. What is the length of the fence?

OpenStudy (anonymous):

Circumference of a circle: \[C=2*\pi*r\] d=2*r

OpenStudy (anonymous):

r=16 yd / 2 = 8 yd C=2*3.1415*8 yd = 50.264 yd